# MMTU: A Massive Multi-Task Table Understanding and Reasoning Benchmark

### v3

@@ A.2 License for existing assets
We provide comprehensive references to the sources of our data assets throughout this paper, with more details in [code-data]. We respect and conform to the licensing terms of existing datasets, whenever we can find such terms when we create our benchmark.
@@ B.1 Column transformation
Program Transform by Example (PTBE) [ptbe-tde, ptbe-transform-prose, transform-column-devlin2017robustfill, transform-column-singh2016blinkfill, he2018transform]. This is a popular task studied in the data management and programming language literature. In this task, we are given a few example output values in a new output column (usually provided by users as demonstrations), and the task is to synthesize a transformation program using input tables, such that the produced output can match the user-given output examples. Figure 1 shows an example of this task. We use the benchmark datasets from [ptbe-tde, ptbe-transform-prose] for this task in MMTU.
Semantic Transform by example (STBE) [sj-sema-join, sj-dataxformer, semantic-join-dtt]. This task is similar to PFBE above, in that users also provide a few example output to demonstrate the intended transformations, but the target transformation requires “semantic” transformations (e.g., country to capital-city, or company to stock-ticker), that cannot be programmed using a syntactic program like in PFBE. We use the datasets from [sj-sema-join, sj-dataxformer]for our benchmark.
Formula by Context (FBC) [fbc-auto-formula, chen2021spreadsheetcoder]. This task is studied in the context of spreadsheets, where given a target spreadsheet cell in which users want to enter a spreadsheet formula, we are asked to predict the intended formula in the target cell. We use the 4 benchmark datasets in [fbc-auto-formula].
@@ B.2 Table Transformation
Table Transform by Output Table (TTBT) [ttbt-autopandas, ttbt-scythe, transform-table-qbo, table-transform-sql-by-output]. In this task, we are given a pair of input/output tables, and the task is to infer a transformation program (in SQL or Pandas), which when executed on the input table, can generate the desired output table. Figure 1 shows an example of this task. We use the datasets from [ttbt-autopandas, ttbt-scythe].
Table Transform by Output Schema (TTBS) [tbps-auto-pipeline, sharma2023automatic, sharma2025datamorpher]. This task is similar to TTBT above, except that the output table is a schematic depiction of how the desired output table should look like, without being the exact output that the desired transformation program would generate on the given input (as it is sometimes hard to generate such output table without first producing the desired program). We use the dataset in [tbps-auto-pipeline]for this task.
Table Transform by Relationalization (TTBR) [ttbr-auto-tables, transform-table-barowy2015flashrelate, transform-table-foofah, xiong2025tablecanoniser, autoprep]. Because relational data analysis often requires input tables to be in a relational form, this task transforms data in non-relational semi-structured forms, into the standard relational form. The task is to predict such a relationalization transformation program, based on the characteristics of the input table. We use the dataset in [ttbr-auto-tables].
@@ B.3 Table matching
Entity Matching (EM) [em-magellandata, em-deepmatcher, em-auto-em, em-ditto, em-book]. Entity matching is the task of determining whether rows or entities from two tables correspond to the same real-world entity. We use the datasets in  [em-magellandata, em-deepmatcher]for benchmarking.
Schema Matching (SM) [sm-valentine, sm-smurf, sm-other-1, schema-matching-cupid]. Schema matching tries to match relevant columns from two tables that map to the same semantic concept. Figure 1 shows an example of this task. We use datasets in [sm-valentine, sm-smurf]for MMTU.
Header Value Matching (HVM) [table-gpt, schema-mapping-survey]. In HVM, we are given a table without headers, and a shuffled list of the original headers in the table, where the task is to match column headers with column values in tables. We use the dataset in [table-gpt]for benchmarking.
@@ B.4 Data cleaning
Data Imputation (DI) [webtables, tablib, table-gpt]. Data imputation is the intuitive task of predicting missing values in a relational table, based on the surrounding context of the table. Figure 1 shows an example of this task. We use the dataset from [webtables, tablib].
Error Detection (ED) [error-detection-survey, error-detect-uni-detect, error-detect-baran, error-detect-auto-test]. The task of Error detection aims to identify erroneous values in a table that are semantically inconsistent or anomalous with the rest of the column. We use the dataset in [error-detect-auto-test].
List to Tables (L2T) [l2t-tegra, list-extraction-2, list-extraction-google, list-other-1]. In List-to-table, the task is to segment records of data without clear separators, into columns, so that values in the same column are homogeneous, and the resulting table becomes a natural relational table, with values consistently aligned in homogenous columns. We use the dataset from [l2t-tegra].
@@ B.5 Table join
Equal Join (EJ) [ej-auto-bi, ej-chen2014fast, ej-fk-ml, ej-hpi]. In Equal-join, we are given a collection of related tables, and the task is to identify all join relationships between the tables. Figure 1 shows an example of this task. We use the dataset in [ej-auto-bi].
Semantic Join (SJ) [sj-sema-join, sj-dataxformer, semantic-join-dtt, auto-map]. In Semantic Join, we are also tasked to join related tables together, but instead of the common equi-join, which is based on string-equality comparisons, the join relationship is based on semantic relatedness (e.g., country and capital city, or company and stock ticket). We use the same semantic-transformation datasets from [sj-sema-join, sj-dataxformer], but converting the input/output transformation columns, as join keys from two separate tables.
@@ B.6 Column relationships
Arithmetic Relationship (AR) [explain-relationship, Auto-relate]. This task focuses on identifying arithmetic relationships from an underlying table. Figure 1 shows an example of this task.
String Relationship (SR) [Auto-relate, ptbe-tde, transform-col-flashfill]. This task focuses on identifying string transformation relationships from a given table.
Functional Relationship (FR) [Auto-relate, fd-1, fd-2]. This task focuses on identifying functional-relationships from input tables. We use the datasets from [Auto-relate] our as benchmarks for all three tasks.
@@ B.7 Table understanding
Needle-in-a-haystack-table (NIHT). In this task, we create a variant of the popular “Needle-in-a-haystack” task in the context of tables, like described in detail in Appendix E. We use the dataset from [tablib]to construct tests in this task.
Needle-in-a-haystack-index (NIHI). In this task, we reverse the NIHT task above, and ask models to identify index positions corresponding to a given value in a table. We use the same dataset from [tablib]to construct tests for this task.
@@ B.8 NL-2-Code
NL-2-SQL (NS) [ns-archer, ns-bird, ns-kaggledbqa, ns-spider, ns-wikisql]. NL-2-SQL is a popular task to translate natural-language questions into SQL statements that can execute given an input table. We use the benchmarks from [ns-archer, ns-bird, ns-kaggledbqa, ns-spider, ns-wikisql]in MMTU.
@@ B.9 Table Question Answering (Table QA)
Table-QA (TQA) [tqa-tablebench, tqa-wikiqa, tqa-finqa, tqa-hitab]. Table QA is another popular task, where models are used to directly answer questions on a given input table, without code execution. We use benchmarks from [tqa-tabfact, tqa-tablebench, tqa-wikiqa, tqa-finqa]for this task.
Fact verification (FV) [tqa-tabfact, tablefact-2, tablefact-3]. Table fact verification is a variant of TQA, in which models are asked if a statement is refuted or supported by facts presented in an input table. We use data from [tqa-tabfact]for this task.
@@ B.10 Knowledge-base mapping (KB Mapping)
Column type annotation (CTA) [semtab19, cta-autotype, cta-doduo, cta-sato, cta-sherlock]. This is a popular task where columns in a table is mapped to known entity types in a knowledge-base. We use the CTA dataset from [semtab19].
Column property annotation (CPA) [semtab19, cpa-2, cpa-3]. [semtab19]. The CPA task is to predict the relationship of a given pair of columns, based on known properties in a knowledge-base. We use the CPA dataset from [semtab19].
Cell entity annotation (CEA) [semtab19, turl, cea-1, cea-2, cea-3]. The CEA task predicts the knowledge-base entity id of a given cell in a table. We use the CEA dataset from [semtab19].
@@ Appendix E Long-context table understanding: A case study of “needle in a haystack in tables”
Figure 10: The needle-in-a-haystack test used in NLP context. Each cell in the heatmap corresponds to a test to retrieve a simple fact planted in a document, for a given document length and retrieval depth. Frontier models like GPT-4o and Gemini-2.5 can now typically achieve perfect accuracy, as shown by the all-green heatmap (results from [niah-2]).
@@
Recall that “Needle-in-a-haystack (NIH)” [NIAH-orig, niah-2], is a popular test traditionally used to evaluate long-context LLM’s ability to retrieve a simple fact (a needle) from a long document context (e.g., 100K or 1M tokens). Recent advances in frontier long-context LLMs have made this task almost irrelevant, as most frontier models such as GPT-4o, Llama-4, and Gemini-2.5 can now score perfectly on NIH [llm-deepseek-v3, gemini-niah, llama-4-release, claude-3-release], like the heatmap in Figure 11 would show. Here the all-green heatmap indicates that this particular LLM under test, GPT-4o, can perfectly retrieve a needle planted at varying depth (y-axis), within documents of varying lengths (x-axis), with 100% accuracy.
@@
The results of NIHT are illustrated also using a heatmap in Figure 11. As we can see, compared to Figure 11, today’s frontier LLMs still face substantial challenges in correctly identifying a cell value within a two-dimensional table, like indicated by the red cells in the heatmap. This is especially true when we increase the number of columns (e.g., with more than 25 columns, LLM accuracy quickly falls below 0.5). Furthermore, we observe a strong asymmetry in the heatmap – e.g., with 50 rows and 5 columns we still have a reasonable accuracy of 0.81 (at the lower-left corner of the heatmap), but with 5 rows and 50 columns the accuracy drops to a lowly 0.43 (at the top-right corner), underscoring the challenge LLMs face when reading “vertically” in the column direction, which are consistent with our intuition that LLMs pre-trained on one-dimensional natural-language texts are less effective in reading two-dimensional tables “vertically”, like was also observed in [table-gpt, table-llm].

### v4

@@ A.2 License for existing assets
We provide comprehensive references to the sources of our data assets throughout this paper, with more details in [95]. We respect and conform to the licensing terms of existing datasets, whenever we can find such terms when we create our benchmark.
@@ B.1 Column transformation
Program Transform by Example (PTBE) [56, 52, 43, 113, 58]. This is a popular task studied in the data management and programming language literature. In this task, we are given a few example output values in a new output column (usually provided by users as demonstrations), and the task is to synthesize a transformation program using input tables, such that the produced output can match the user-given output examples. Figure 1 shows an example of this task. We use the benchmark datasets from [56, 52] for this task in MMTU.
Semantic Transform by example (STBE) [57, 1, 37]. This task is similar to PFBE above, in that users also provide a few example output to demonstrate the intended transformations, but the target transformation requires “semantic” transformations (e.g., country to capital-city, or company to stock-ticker), that cannot be programmed using a syntactic program like in PFBE. We use the datasets from [57, 1]for our benchmark.
Formula by Context (FBC) [23, 25]. This task is studied in the context of spreadsheets, where given a target spreadsheet cell in which users want to enter a spreadsheet formula, we are asked to predict the intended formula in the target cell. We use the 4 benchmark datasets in [23].
@@ B.2 Table Transformation
Table Transform by Output Table (TTBT) [11, 125, 122, 142]. In this task, we are given a pair of input/output tables, and the task is to infer a transformation program (in SQL or Pandas), which when executed on the input table, can generate the desired output table. Figure 1 shows an example of this task. We use the datasets from [11, 125].
Table Transform by Output Schema (TTBS) [135, 110, 111]. This task is similar to TTBT above, except that the output table is a schematic depiction of how the desired output table should look like, without being the exact output that the desired transformation program would generate on the given input (as it is sometimes hard to generate such output table without first producing the desired program). We use the dataset in [135]for this task.
Table Transform by Relationalization (TTBR) [78, 10, 67, 132, 72]. Because relational data analysis often requires input tables to be in a relational form, this task transforms data in non-relational semi-structured forms, into the standard relational form. The task is to predict such a relationalization transformation program, based on the characteristics of the input table. We use the dataset in [78].
@@ B.3 Table matching
Entity Matching (EM) [38, 98, 146, 80, 102]. Entity matching is the task of determining whether rows or entities from two tables correspond to the same real-world entity. We use the datasets in  [38, 98]for benchmarking.
Schema Matching (SM) [71, 145, 12, 89]. Schema matching tries to match relevant columns from two tables that map to the same semantic concept. Figure 1 shows an example of this task. We use datasets in [71, 145]for MMTU.
Header Value Matching (HVM) [79, 105]. In HVM, we are given a table without headers, and a shuffled list of the original headers in the table, where the task is to match column headers with column values in tables. We use the dataset in [79]for benchmarking.
@@ B.4 Data cleaning
Data Imputation (DI) [16, 44, 79]. Data imputation is the intuitive task of predicting missing values in a relational table, based on the surrounding context of the table. Figure 1 shows an example of this task. We use the dataset from [16, 44].
Error Detection (ED) [31, 126, 90, 21]. The task of Error detection aims to identify erroneous values in a table that are semantically inconsistent or anomalous with the rest of the column. We use the dataset in [21].
List to Tables (L2T) [30, 17, 45, 75]. In List-to-table, the task is to segment records of data without clear separators, into columns, so that values in the same column are homogeneous, and the resulting table becomes a natural relational table, with values consistently aligned in homogenous columns. We use the dataset from [30].
@@ B.5 Table join
Equal Join (EJ) [81, 27, 108, 65]. In Equal-join, we are given a collection of related tables, and the task is to identify all join relationships between the tables. Figure 1 shows an example of this task. We use the dataset in [81].
Semantic Join (SJ) [57, 1, 37, 128]. In Semantic Join, we are also tasked to join related tables together, but instead of the common equi-join, which is based on string-equality comparisons, the join relationship is based on semantic relatedness (e.g., country and capital city, or company and stock ticket). We use the same semantic-transformation datasets from [57, 1], but converting the input/output transformation columns, as join keys from two separate tables.
@@ B.6 Column relationships
Arithmetic Relationship (AR) [112, 7]. This task focuses on identifying arithmetic relationships from an underlying table. Figure 1 shows an example of this task.
String Relationship (SR) [7, 56, 51]. This task focuses on identifying string transformation relationships from a given table.
Functional Relationship (FR) [7, 103, 129]. This task focuses on identifying functional-relationships from input tables. We use the datasets from [7] our as benchmarks for all three tasks.
@@ B.7 Table understanding
Needle-in-a-haystack-table (NIHT). In this task, we create a variant of the popular “Needle-in-a-haystack” task in the context of tables, like described in detail in Appendix E. We use the dataset from [44]to construct tests in this task.
Needle-in-a-haystack-index (NIHI). In this task, we reverse the NIHT task above, and ask models to identify index positions corresponding to a given value in a table. We use the same dataset from [44]to construct tests for this task.
@@ B.8 NL-2-Code
NL-2-SQL (NS) [147, 77, 74, 138, 148]. NL-2-SQL is a popular task to translate natural-language questions into SQL statements that can execute given an input table. We use the benchmarks from [147, 77, 74, 138, 148]in MMTU.
@@ B.9 Table Question Answering (Table QA)
Table-QA (TQA) [131, 137, 28, 29]. Table QA is another popular task, where models are used to directly answer questions on a given input table, without code execution. We use benchmarks from [24, 131, 137, 28]for this task.
Fact verification (FV) [24, 136, 141]. Table fact verification is a variant of TQA, in which models are asked if a statement is refuted or supported by facts presented in an input table. We use data from [24]for this task.
@@ B.10 Knowledge-base mapping (KB Mapping)
Column type annotation (CTA) [66, 134, 116, 140, 61]. This is a popular task where columns in a table is mapped to known entity types in a knowledge-base. We use the CTA dataset from [66].
Column property annotation (CPA) [66, 36, 99]. [66]. The CPA task is to predict the relationship of a given pair of columns, based on known properties in a knowledge-base. We use the CPA dataset from [66].
Cell entity annotation (CEA) [66, 42, 13, 130, 83]. The CEA task predicts the knowledge-base entity id of a given cell in a table. We use the CEA dataset from [66].
@@ Appendix E Long-context table understanding: A case study of “needle in a haystack in tables”
Figure 10: The needle-in-a-haystack test used in NLP context. Each cell in the heatmap corresponds to a test to retrieve a simple fact planted in a document, for a given document length and retrieval depth. Frontier models like GPT-4o and Gemini-2.5 can now typically achieve perfect accuracy, as shown by the all-green heatmap (results from [107]).
@@
Recall that “Needle-in-a-haystack (NIH)” [87, 107], is a popular test traditionally used to evaluate long-context LLM’s ability to retrieve a simple fact (a needle) from a long document context (e.g., 100K or 1M tokens). Recent advances in frontier long-context LLMs have made this task almost irrelevant, as most frontier models such as GPT-4o, Llama-4, and Gemini-2.5 can now score perfectly on NIH [82, 119, 120, 32], like the heatmap in Figure 11 would show. Here the all-green heatmap indicates that this particular LLM under test, GPT-4o, can perfectly retrieve a needle planted at varying depth (y-axis), within documents of varying lengths (x-axis), with 100% accuracy.
@@
The results of NIHT are illustrated also using a heatmap in Figure 11. As we can see, compared to Figure 11, today’s frontier LLMs still face substantial challenges in correctly identifying a cell value within a two-dimensional table, like indicated by the red cells in the heatmap. This is especially true when we increase the number of columns (e.g., with more than 25 columns, LLM accuracy quickly falls below 0.5). Furthermore, we observe a strong asymmetry in the heatmap – e.g., with 50 rows and 5 columns we still have a reasonable accuracy of 0.81 (at the lower-left corner of the heatmap), but with 5 rows and 50 columns the accuracy drops to a lowly 0.43 (at the top-right corner), underscoring the challenge LLMs face when reading “vertically” in the column direction, which are consistent with our intuition that LLMs pre-trained on one-dimensional natural-language texts are less effective in reading two-dimensional tables “vertically”, like was also observed in [79, 117].